OpenStudy (anonymous):

5/6x + 2/3 = -1/6x - 5/3

OpenStudy (anonymous):

\[\frac{5}{6x}+\frac{2}{3}=\frac{1}{6x}-\frac{5}{3}\]?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[\frac{5}{6x}+\frac{2}{3}=\frac{1}{6x}-\frac{5}{3}\]\[\frac{4}{6x}=-\frac{7}{3}\]\[-42x=12\]\[x=-\frac{12}{42}=\frac{2}{7}\]

OpenStudy (anonymous):

denominator is 6x\[4x+10x=-1-5\]\[5+(2*2x)=-1-(5*2x)\]\[5+4x=1-10x\]\[4x+10x=-1-5\]\[14x=-6\]\[x=-\frac{6}{14}\]\[x=-\frac{3}{7}\]

OpenStudy (anonymous):

I mean -2/7.

OpenStudy (anonymous):

carefully the sign i first forget forgot -1/6x not positive 1/6x

OpenStudy (anonymous):

yeah, I solved the one you wrote, Nancy. oops!